Tried several times and kept getting the wrong answer.
media%2F664%2F6649e78e-2a6b-469a-a6ce-a9
Answer #1

Q = CV
V = Q/C = 25µ/15µ = 1.67 V initial
final, V = 10/15 = 0.67 V

voltage on a cap, discharging
V = V0exp(–t/?)

RC = ? = time constant

? = 15µ x 1.0k = 15 ms
V = V0exp(–t/?)
0.67 = 1.67exp(–t/15m)
exp(–t/15m) = 0.401
–t/15m = -0.9133
t = (15ms)(0.9133) = 13.69 ms  =0.0137 s